A random sample of 25 with a mean of 80 and a standard deviation of 30 is taken from a population of 1000 that is normally distributed. Find a) the 90% b) the 95% and c) the 99% confidence intervals for the unknown population mean.
1
Expert's answer
2019-07-31T09:19:31-0400

Let  α1 = 1 - 0.90 = 0.1. α2 = 1-0.95=0.05, α3 = 1 - 0.99=0.01 are significance levels, degree of freedom

d = 25 - 1 = 24.

For α1 = 0.1 and d = 24 critical t-value t1 = 1.711.

For α2 = 0.05 and d = 24 critical t-value t2 = 2.064,

For α3 = 0.1 and d = 24 critical t-value t3 = 2.797.

We will find corresponding confidence intervals, using expression


"(\mu-\frac{t\cdot s}{\sqrt{n}}, \mu+\frac{t\cdot s}{\sqrt{n}})"

"where\ \mu\ is \ sample\ mean ,\ t\ is\ a\ critical\ value,"

"n\ is\ the\ size\ and\ s\ is\ standard\ deviation\ of\ the\ sample."

a) Confidence interval is


"(80 - \frac{1.711\cdot 30}{\sqrt{25}}, 80 + \frac{1.711\cdot 30}{\sqrt{25}})="

"(80-10.266,\ 80+10.266)="

"(69.734,\ 90.266)"

b) Confidence interval is


"(80 - \frac{2.064\cdot 30}{\sqrt{25}}, 80 + \frac{2.064\cdot 30}{\sqrt{25}})="

"(80-12.384,\ 80+12.384)="

"(67.616,\ 92.384)"

c) Confidence interval is


"(80 - \frac{2.797\cdot 30}{\sqrt{25}}, 80 + \frac{2.797\cdot 30}{\sqrt{25}})="

"(80-16.782,\ 80+16.782)="

"(63.218,\ 96.782)"

Answer: a) (69.734, 90.266), b) (67.616, 92.384), c) (63.218, 96.782)